• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name X. Zheng

    Log entry time 16:11:08 on March 08, 2012

    Entry number 364662

    keyword=Day shift summary

    08:00 Left HRS is coming back to 2.251 GeV/c;
    08:10 Left HRS momentum is set;
    08:15 Escorted controlled access, Jixie is inserting the sieve slits to
    both HRSs;
    08:36 End of access but MCC is having problems to deliver the beam;
    09:05 Septum current is ramped to 725.40 Amp;
    09:50 MCC is ready to send beam. Target experts will do the target
    position calibration first with the "empty hole" aluminum target.
    09:58 50nA, fast raster 2x2, slow raster 1.4V 
    10:22 Beam is down again and Hall C is using this time to change their HWP;
    12:57 beam is back!  We continue with the target position calibration.
    Fast raster is OFF now and slow raster is ON;
    14:55 We finished all target position calibration and we are now starting
    optics. It took a while both for us to convey the correct beam nominal
    position to the MCC, and for the MCC to get to these requested positions
    BPMAX: -0.3 (H05AX)
    BPMAY: -1.4 (H05AY)
    BPMBX: -1.1 (H05BX)
    BPMBY: -2.7 (H05BY)
    15:30 MCC called and said they will take the beam away again for 15
    minutes to add a patch to the gun.
    15:55 Beam off for the gun patch, end runs, LHe are not fully drained yet.
    Run list for checking target position using the "empty hole" target.
    50nA, fast raster 2x2, slow raster 1.4V:
    Optics Run list:
    With LHe, Pointing, Sieve in, 40mil C12 target: 
    2919, 2920, 2921, 2922
    22015, 22016, 22017, 22018
    Draining LHe, Pointing, Sieve in, 40mil C12 target:
    (LHe was not fully drained at the end of run due to MCC test)